OpenStudy (anonymous):

There are 96new houses being built in a neighborhood. Last month, 1/3 of them were sold. This month, 3/8 of the remaining houses were sold. How many houses are left to be sold?

OpenStudy (anonymous):

Sixty?

OpenStudy (anonymous):

how did u get that ?

OpenStudy (anonymous):

I subtratced a third and three eighths from 96.

OpenStudy (anonymous):

subtracted*

OpenStudy (anonymous):

WAIT NO.

OpenStudy (anonymous):

The answer is 28 houses.

OpenStudy (anonymous):

r u sure?

OpenStudy (anonymous):

Yes. 3/8 = 36. How do I know? 96/8=12, and 12*3=36. Also, a third of 96 is 32, so you would subtract 32 from 96 which is 64, and then you'd subtract 36 more. 36 came from 3/8.

OpenStudy (anonymous):

it said that was wrong the answer was 40

OpenStudy (anonymous):

Fml.